# Tracking Meets LoRA: Faster Training, Larger Model, Stronger Performance

### v1

@@ 4.3 Ablation Experiments
Table 4: Ablation comparison of LoRA vs. other PEFT methods. The base model is LoRAT-L-224 (last line). $\mathrm{LoRA}_{qv}$ indicates the LoRA module is added to the query and value projection matrix of the ViT backbone. $\mathrm{LoRA}_{all}$ indicates the LoRA module is added to all linear layers of the ViT backbone.
@@
- 73.5 79.8 50.3 57.3 59.3 61.4 307 100 119
Adapter [50] 60.4 61.1 41.9 46.2 50.3 49.3 314 3.3 78
VPT-Deep ($p=5$[32] 65.8 68.7 45.6 51.6 53.9 53.9 308 1.4 83
VPT-Deep ($p=10$[32] 66.2 69.5 47.1 53.5 54.7 55.1 308 1.5 83
$\mathrm{(IA)}^{3}$ [42] 58.9 59.0 39.8 43.5 47.1 48.0 308 1.4 119
$\mathrm{LoRA}_{qv}(r=4)$ [30] 66.2 69.8 46.8 53.1 55.2 54.6 308 1.5 119
$\mathrm{LoRA}_{all}(r=16)$ [30] 72.9 79.0 50.9 58.2 58.9 61.0 315 3.6 119
$\mathrm{LoRA}_{all}(r=64)$ [30] 74.3 80.8 51.5 58.7 59.4 61.9 336 9.7 119
Input Embedding. This ablation verifies the effectiveness of the proposed decoupled input embedding scheme. The experiments are conducted on the ViT-L [15] backbone with two input resolution settings. From the results in Table 5, we observe that: 1) The proposed input embedding scheme achieves the best results in both resolution settings with LoRA enabled. 2) The OSTrack-like positional embedding scheme performs best on the LaSOT benchmark when the model is trained with full fine-tuning. However, it lags behind our proposed input embedding solution when the model is trained with LoRA. 3) For multi-resolution adaptation of absolute positional embedding, slicing-based adaptation is better than interpolation-based adaptation.
Table 5: Ablation on the choice of input embeddings. All experiments are conducted on ViT-L [15] backbone with two input resolution settings. ❶: sinusoidal position encoding, MixViT [9] like; ❷: a pair of absolute positional embeddings interpolated from the pre-trained ViT models, OSTrack [62] like; ❸: Ours, shared absolute positional embedding with token type embedding; ❹: Ours, ❷ with token type embedding and foreground object indicator. ❺: Ours, shared absolute positional embedding with token type embedding and foreground object indicator. The positional embeddings in the variants with LoRA enabled are frozen during training.
@@
SUC P SUC P SUC P
224 69.8 74.4 47.9 54.4 55.2 55.9
224 73.7 80.1 50.3 57.5 58.7 60.8
224 73.5 79.8 50.3 57.3 59.3 61.4
224 73.1 74.4 47.9 54.4 55.2 55.9
224 73.4 80.1 51.2 58.8 58.9 61.1
224 74.1 80.7 51.6 59.1 59.3 61.7
224 73.7 80.2 50.6 57.7 58.9 61.2
224 74.3 82.2 51.5 58.7 59.4 61.9
378 74.0 80.6 52.8 60.5 59.9 62.3
378 74.6 81.6 52.9 60.6 60.2 63.1
378 75.1 82.2 53.1 60.8 60.7 63.7
378 75.4 82.2 53.5 61.4 60.8 63.9
Table 6: Ablation on network modules. Conv head: Replace the 3-layers MLP-base head in our tracker with 5 stacked Conv-BN-ReLU layers; OSTrack [62] head: Replace the head in our tracker with OSTrack [62] convolution-based head. S-MAM: Replace the self-attention module in ViT backbone with Slimming Mixed Attention Module [9].
[…]
LoRA L-224 74.3 83.6 80.8 51.5 65.0 58.7 59.4 61.9
+ S-MAM [9] 73.2 82.4 79.9 50.5 63.7 57.4 59.3 61.5
+ Conv head 4.7 6.4 1.6 4.5 9.4 0.8 0.0 0.1
+ OSTrack [62] head 5.4 5.1 1.3 4.9 9.0 1.2 0.1 0.1
Full fine-tune L-224 73.5 82.4 79.8 50.3 63.2 57.3 59.3 61.4
+ S-MAM [9] 73.2 82.2 79.6 50.1 63.2 57.4 58.3 60.5
+ Conv head 72.8 81.7 79.0 51.1 64.2 58.4 58.1 60.1
+ OSTrack [62] head 73.7 82.4 79.9 48.1 60.4 54.2 59.4 61.5
Full self-attention vs. S-MAM. Slimming Mixed Attention Module (S-MAM) is a variant of self-attention proposed in MixViT [9], pruning the template-to-search area cross-attention. We compare the performance of full self-attention (performing self-attention operation on flattened template and search region tokens) and S-MAM in Table 6. Full self-attention outperforms S-MAM in both fine-tuning with LoRA and full fine-tuning settings. The performance of our tracker with S-MAM also gets better result on LoRA-based fine-tuning, verifying that our model design can adapt to other one-stream tracking models.
MLP-only head network. We evaluate the performance of MLP-only and convolutional head networks under different fine-tuning settings. As shown in Table 6, the OSTrack [62] convolutional head outperforms the MLP-only head on the full fine-tuning setting but fails to converge on the LoRA-based fine-tuning setting. We believe this is due to the inductive biases in the convolution layer, which may not meet the low "intrinsic dimension" assumption required by low-rank adaptation. Additionally, an MLP-only head is computationally efficient.
Aligned Comparison with OSTrack. We provide aligned performance comparison with OSTrack [62] by replacing only the backbone. OSTrack is trained without the early candidate elimination module for fair comparison. Input resolution of variants is adjusted according to the ViT patch size. As shown in Table 7, all variants of OSTrack are inferior to our base model. Besides, the performance of the DINOv2 backbone on OSTrack lags behind the default setting (MAE backbone). The results demonstrate the effectiveness of our model design.
[…]
Different Pre-training Methods. We conduct experiments on four ViT variants with different pre-training methods [26, 51, 19, 49]. From Table 9, we observe that the self-supervised DINOv2 outperforms all other ViT variants on most datasets. Compared with the results in Table 7, our model design can correctly unleash the power of large-scale pre-training.
Training efficiency of LoRA. We verify the training efficiency of LoRA on two machines, one with 8 NVIDIA V100 GPUs (Table 8), one with a single NVIDIA 4090 GPU (Table 10). The results show that fine-tuning with LoRA significantly saves training time and GPU memory, especially on large models.
Due to space limitation, please refer to supplementary material for more experimental results and analysis.
5 Conclusion
In this work, we introduce a novel approach to visual tracking by applying Low Rank Adaptation (LoRA) within a one-stream tracking framework. Our contributions include two innovative designs that enable effective adaptation of pre-trained models for tracking, significantly reducing the resource requirements. The proposed tracker, LoRAT, not only achieves state-of-the-art performance on multiple benchmarks but also demonstrates the feasibility of training advanced tracking models with manageable resources. This study bridges the gap between advanced tracking technologies and resource accessibility, marking a pivotal step towards sustainable advancements in the field of visual tracking.

### v2

@@ 0.C.3 Performance under Different Initialization Methods of LoRA Weights
Table 15: Ablation comparison of LoRA vs. other PEFT methods. The base model is LoRAT-L-224 (last line).
@@
- 73.0 79.3 52.6 59.8 60.8 64.3 307 100 119
Adapter [55] 55.0 52.9 38.3 39.8 45.4 39.2 314 3.3 78
VPT-Deep ($p=5$[33] 58.0 56.8 41.3 44.2 48.0 43.5 308 1.4 83
VPT-Deep ($p=10$[33] 59.3 58.9 42.1 45.4 49.4 45.5 308 1.5 83
$\mathrm{(IA)}^{3}$ [45] 59.5 60.0 43.1 47.5 49.4 47.0 308 1.4 119
$\mathrm{LoRA}_{qv}(r=4)$ 66.2 70.0 48.5 54.9 55.7 57.1 308 1.5 119
$\mathrm{LoRA}_{all}(r=16)$ 73.0 79.3 52.9 60.4 60.4 64.3 315 3.6 119
$\mathrm{LoRA}_{qv}(r=64)$ 71.9 77.7 52.9 59.2 59.4 63.0 314 3.3 119
$\mathrm{LoRA}_{qkvo}(r=64)$ 72.9 79.5 52.7 60.3 60.5 64.7 320 5.3 119
$\mathrm{LoRA}_{mlp}(r=64)$ 72.8 78.9 52.2 59.6 60.3 63.9 323 6.2 119
$\mathrm{LoRA}_{all}(r=64)$ 74.2 80.9 52.8 60.0 61.1 65.1 336 9.7 119
0.C.4 Comparison with other PEFT methods
In this section, we evaluate the effectiveness of Low-Rank Adaptation (LoRA) compared to other PEFT methods within our tracking framework. We conduct a comprehensive comparison using three representative PEFT approaches: Adapter [30], a representative of adatper-based PEFT method [30, 44, 55]. We follow the original design as in [30], insert the adapter layers between the multi-head self-attention module and the subsequent residual connection; Visual Prompt Tuning (VPT) [33], a representative of prompt tuning-based PEFT method for vision model; $\mathrm{(IA)}^{3}$ [45], a recent zero inference computational cost PEFT method. Furthermore, we explore various configurations of LoRA, designated as: $\mathrm{LoRA}_{qv}$: LoRA added to query and value projection matrix of ViT backbone. $\mathrm{LoRA}_{qkvo}$: LoRA added to all linear layers of multi-head self-attention modules of ViT backbone. $\mathrm{LoRA}_{mlp}$: LoRA added to all linear layers of MLP layers of ViT backbone. $\mathrm{LoRA}_{all}$: LoRA added to every linear layers of ViT backbone.
Using LoRAT-L-224 as the base model, results from Table 15 illustrate that LoRA significantly outperforms other PEFT methods under various settings. Notably, $\mathrm{LoRA}_{qv}(r=4)$ surpasses all compared PEFT methods with only 1.5% trainable parameters. $\mathrm{LoRA}_{all}(r=16)$ matches the performance of full fine-tuning with 3.6% trainable parameters. Additionally, from the bottom section of Table 15, we observe that applying LoRA to both MHSA and MLP obtains the best performance.
Table 16: Ablation on attention mechanism. S-MAM: Replace the self-attention module in ViT backbone with Slimming Mixed Attention Module [10].
@@ 0.C.4 Comparison with other PEFT methods
SUC PNorm P SUC PNorm P SUC P
LoRA:
L-224 74.2 83.6 80.9 52.8 64.7 60.0 61.1 65.1
+ S-MAM [10] 73.5 82.6 80.1 52.4 64.1 59.8 61.0 64.9
Full Fine-tune:
L-224 73.0 81.8 79.3 52.6 64.1 59.8 60.8 64.3
+ S-MAM [10] 73.0 82.1 79.6 52.5 64.0 59.9 60.0 63.2
0.C.5 Full self-attention vs. S-MAM
Slimming Mixed Attention Module (S-MAM) is a variant of self-attention introduced in MixViT [10], which specifically prunes the template-to-search area cross-attention. We evaluate and compare the performance of full self-attention (performing the self-attention operation on flattened template and search region tokens) and S-MAM in Table 16. Our findings indicate that full self-attention consistently outperforms S-MAM in both LoRA-based and full fine-tuning settings. Despite this, our tracker with S-MAM still shows improved results under LoRA-based fine-tuning compared to full fine-tuning. This demonstrates the adaptability of our model design to various one-stream tracking architectures.